An 8.5% coupon bond pays interest semiannually and has 11.5 years to maturity. The bond has a face value of $1,000 and a market value of $878.50. What is the yield to maturity? Is it A.11.5% , B.6.75%, C.10.33%, D. 8.78%
Answer #1

Number or periods = 11.5 * 2 = 23

Coupon = (0.085 * 1000) /2 = 42.5

Yield to maturity = 10.33%

Keys to use in a financial calculator:

2nd I/Y 2

FV 1000

PV -878.50

PMT 42.5

N 23

CPT I/Y